Please help me solve this equation:
The sum of four positive integers is 68. The product of the same four integers is 50141. The sum of the squares of the same four integers is 1420. What are the values of the four positive integers?
a+b+c+d=68
axbxcxd=50141
a^2+b^2+c^2+d^2=1420

Answer by Alan3354(69443)

The sum of four positive integers is 68. The product of the same four integers is 50141. The sum of the squares of the same four integers is 1420. What are the values of the four positive integers?
a+b+c+d=68
axbxcxd=50141
a^2+b^2+c^2+d^2=1420
------------------------
There are 4 variables but only 3 equations. There are an infinite number of solutions, but only one solution with integers.
----------
a*b*c*d=50141 --> all 4 numbers are odd (use * for multiply, not x)
Factor 50141
50141 = 7*13*19*29